OpenStudy (anonymous):

I am having real trouble with linear equation with fractions, my teacher is bollocks.

OpenStudy (saifoo.khan):

\[\Huge y=mx +b\]

OpenStudy (anonymous):

never seen that in my life

OpenStudy (anonymous):

m will be the slope

OpenStudy (saifoo.khan):

LOOL, ok so what are you stuck with?

OpenStudy (anonymous):

b will be the y intercept. :))

OpenStudy (anonymous):

solve for x 1/21=1/2x+x+1/3

OpenStudy (saifoo.khan):

That is as hard as a Diamond.

OpenStudy (anonymous):

really so I am not stupid?

OpenStudy (saifoo.khan):

HAHA.

OpenStudy (saifoo.khan):

SOLVE IT.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

you can do it.

OpenStudy (saifoo.khan):

Righto.

OpenStudy (anonymous):

if I tell my you guys the steps I take maybe you can tell me where I am going wrong

OpenStudy (saifoo.khan):

Great.

OpenStudy (anonymous):

first I subtracted 1/3 from either side

OpenStudy (anonymous):

wait

OpenStudy (saifoo.khan):

http://www.twiddla.com/611673

OpenStudy (anonymous):

just separate the x from the numbers

OpenStudy (anonymous):

put your x on the other side and the numbers on the right

OpenStudy (anonymous):

x = n "n for numbers, real numbers"